OpenStudy (coldplay):

(Points : 1) x ≤ –10 x ≥ –10 x ≤–3 x ≥ –3 Question 2.2. Solve. –5t ≥ 65 (Points : 1) t ≥ 13 t ≤ 13 t ≥ –13 t ≤ –13

OpenStudy (coldplay):



OpenStudy (coldplay):

#1 plz ignore didn;t find the equation

OpenStudy (coldplay):

#2 plz help me with thanks

OpenStudy (anonymous):

divide both sides by -5 Flip the sign cuz u divided with a negative and u get t ≤ –13

OpenStudy (anonymous):

The last one t<-13

OpenStudy (coldplay):

Thank you roz_mar23

OpenStudy (anonymous):

ur welcome

OpenStudy (coldplay):

Ur the best

OpenStudy (anonymous):

Thanx

OpenStudy (coldplay):



OpenStudy (coldplay):

Is subtraction Commutative? Use an example to show your reasoning. I need help

OpenStudy (anonymous):

For example, when we say addition is commutative over the set of real numbers, we mean that a + b = b + a for all real numbers a and b. Subtraction is not commutative over real numbers since we can't say that a – b = b – a for all real numbers a and b.

OpenStudy (anonymous):

do you still need help?

OpenStudy (coldplay):

Yes I do.........